For each of the following sets of pressure/volume data, calculate the new volume of the gas sample after the pressure changeWrite answer to correct number of significant figures.
Answer #1

a)

Given:

Vi = 101 mL

Pi = 704 mmHg

Pf = 711 mmHg

use:

Pi*Vi = Pf*Vf

704 mmHg * 101 mL = 711 mmHg * Vf

Vf = 100. mL

Answer: 100. mL

b)

Given:

Vi = 217 mL

Pi = 1.02 atm

Pf = 0.953 atm

use:

Pi*Vi = Pf*Vf

1.02 atm * 217 mL = 0.953 atm * Vf

Vf = 232 mL

Answer: 232 mL

c)

Given:

Vi = 3.23 L

Pi = 105 kPa

Pf = 107 kPa

use:

Pi*Vi = Pf*Vf

105 kPa * 3.23 L = 107 kPa * Vf

Vf = 3.17 L

Answer: 3.17 L